A 4 x 4 x 4 cube consisting of smaller cubes is painted and then broken apart. How many of the smaller cubes will have exactly 2 painted sides?
(A) 8 (B) 16 (C) 20 (D) 24 (E) 32

Answer by richwmiller(17219)

Your problem did not say how big or small the smaller cubes are. Are they 2by2by2 or 1 by 1 by1 or something in between?
Assuming 1 by1
The extreme corners of the 4 by 4 by 4 will have three sides painted
the top and bottom rows then minus the four corners are 2*4*2=16
but the side edges of the 4by4by4 will also have 2 sides painted.
2*4=8
16+8=24 painted on 2 sides
D) 24